OpenStudy (anonymous):

\[\sqrt[3]{x}\] derived

OpenStudy (anonymous):

take the derivative?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\(\large \sqrt[3]{x} = x^{\frac{1}{3}}\) use the power rule...

OpenStudy (anonymous):

\(\large \frac{d}{dx}x^{\frac{1}{3}}=\frac{1}{3}x^{-\frac{2}{3}} \)